OpenStudy (anonymous):

Determine the mass of a shape over [0,1]x[5,8]x[9,10]. with density at anypoint given by density function f(x,y,z) = xyz .

OpenStudy (bobbyleary):

\[M = \int\limits_{0}^{1}\int\limits_{5}^{8}\int\limits_{9}^{10} (xyz)dzdydx\] For other problems, if a density function is not giving, you are to assume the object has an evenly distributed mass so f(x,y,x) = 1. Which means: \[M = \int\limits_{}^{}\int\limits_{}^{}\int\limits_{}^{}(1)dxdydz\]

OpenStudy (bobbyleary):

Let me know if you have issues solving the triple integral.

OpenStudy (anonymous):

Thanks man! :)